How long will it take for a 40.0 gram sample of I-131 (half life = 8.040 days) to decay to 1/100 its original mass?
Expert's answer
1
Expert's answer
2020-10-30T06:43:34-0400

The general formula is

"t=\frac{t_{1/2}ln(\frac{N_t}{N_0})}{-ln2}" , where "N_t" is the amount of the substance at given time, "N_0" is the initial amount, "t" is the time and "t_{1\/2}" is the half life.

Therefore,

"t=\frac{8.040days\times{ln\frac{1}{100}} }{-0.693}=53.4days"


Answer: 53.4 days